# A Laboratory Experiment on Using Different Financial-Incentivization Schemes in Software-Engineering Experimentation

### v5

@@ Abstract
In software-engineering research, many empirical studies are conducted with open-source or industry developers. However, in contrast to other research communities like economics or psychology, only few experiments use financial incentives (i.e., paying money) as a strategy to motivate participants’ behavior and reward their performance. The most recent version of the SIGSOFT Empirical Standards mentions payouts only for increasing participation in surveys, but not for mimicking real-world motivations and behavior in experiments. Within this article, we report a controlled experiment in which we tackled this gap by studying how different financial incentivization schemes impact developers. For this purpose, we first conducted a survey on financial incentives used in the real-world, based on which we designed three incentivization schemes: (1) a performance-dependent scheme that employees prefer, (2) a scheme that is performance-independent, and (3) a scheme that mimics open-source development. Then, using a between-subject experimental design, we explored how these three schemes impact participants’ performance. Our findings indicate that the different schemes can impact participants’ performance in software-engineering experiments. Due to the small sample sizes, our results are not statistically significant, but we can still observe clear tendencies. Our contributions help understand the impact of financial incentives on participants in experiments as well as real-world scenarios, guiding researchers in designing experiments and organizations in compensating developers.
@@ 1 Motivation
Interestingly, there are no guidelines or recommendations on using financial incentives in software-engineering experimentation. Namely, the current SIGSOFT Empirical Standards1 (Ralph, 2021) (as of August 4, 2021; commit b046f37) mention incentives solely in the context of rewarding participation in surveys. For instance, the current SIGSOFT Empirical Standards [^1] https://github.com/acmsigsoft/EmpiricalStandards (Ralph, 2021), as of January 22, 2024 (commit 9374ea5), mention incentives solely in the context of longitudinal studies and rewarding participation in surveys to increase participation. Also, to the best of our knowledge and based on a literature review, financial incentives that reward participants’ performance during an experiment are not used systematically in empirical software engineering. Although some studies broadly incentivize performance (e.g., Sayagh et al. (2020) or Shargabi et al. (2020)), these do not aim to improve the validity of the experiment, only participation. Furthermore, we know from experimental economics (Charness and Kuhn, 2011; Carpenter and Huet-Vaughn, 2019) that finding a realistic (and thus externally valid) way to reward performance is challenging and no simple one-fits-all solution exists. For instance, the performance of open-source developers depends less on financial rewards than those of industrial developers (Baddoo et al., 2006; Ye and Kishida, 2003; Huang et al., 2021; Beecham et al., 2008).
@@
As a step towards understanding and systematizing the potential of using financial incentives in software engineering experimentation, we propose have conducted a two-part study comprising a survey and a controlled experiment in the context of bug detection through code reviews (Krüger et al., 2022). First, we will conduct used a survey with practitioners to elicit real-world incentivisation schemes on bug finding. In the survey, we will distinguish distinguished between the schemes most participants prefer and those actually employed. Building on the results, we will define two corresponding defined one payoff function for our experiment. Please note that we originally planned to have two functions from the survey, one for the most applied (MA) and one for the most preferred (MP) incentives (Krüger et al., 2022). However, the survey responses for the MA incentives were identical to no performance-based incentives, which we added as a control treatment anyway. To extend our experiment, we will add added two more payoff functions: one that is performance-independent and one that resembles the motives of open-source developers. We derive derived the latter function using the induced-value method established in experimental economics (Smith, 1976; Weimann and Brosig-Koch, 2019), which induces a controlled willingness of participants to achieve a desired goal (i.e., identify a bug) or obtain a certain good during an experiment by mimicking its monetary value (e.g., a reward). Second, we will employ employed our actual between-subject experiment to explore to what extent each of the four three payoff functions impacts the participants’ behavior. Overall, we primarily contribute to improving researchers’ understanding of whether and how financial incentives can help software engineering experimentation. However, our experiment also has the potential to can also help reveal whether different incentivisation schemes could improve practitioners’ motivation. Our survey and experimental design artifacts are available for peer-reviewing.
@@
    We share our artifacts, including the design and results of our survey as well as experiment in anonymous form within a persistent open-access repository. [^2] Will be published on Zenodo
@@ 2 Related Work
Experiments in software engineering are comparable to “real-effort experiments” in experimental economics, which involve participants who solve certain tasks to increase their payoffs. Consequently, we build built on experiences from the field of experimental economics, which involves a large amount of literature on how and when to use financial incentives in real-effort experiments (van Dijk et al., 2001; Greiner et al., 2011; Gill and Prowse, 2012; Erkal et al., 2018). For instance, some findings indicate gender differences regarding the impact of incentivization schemes, which we have to consider during our experiment. In detail, research has shown that men choose more competitive schemes (e.g., tournaments, performance payments). Similarly, participants with higher social preferences select such competitive schemes more rarely (Niederle and Vesterlund, 2007; Dohmen and Falk, 2011). We will consider considered such factors when analyzing the results of our experiment (e.g., comparing gender differences if the number of participants allows).
@@
Software-engineering researchers have investigated the motivations of open-source developers to a much greater extent (Gerosa et al., 2021; Hertel et al., 2003; Hars and Ou, 2002; Ye and Kishida, 2003; Huang et al., 2021). From the economics perspective, open-source systems represent a public good (Bitzer et al., 2007; Lerner and Tirole, 2003): they are available to everyone and their consumption do not yield disadvantages to anyone else. A typical problem of public goods is that individual and group incentives collide, which usually leads to an insufficient provision of the good. While typical explanations for open-source development focus on high intrinsic motivation to contribute or learn, this is not always the case. For instance, Roberts et al. (2006) show that financial incentives can actually improve open-source developers’ motivation (in terms of contributions). Still, financial incentives are at least not always the predominant motivators for software developers (Beecham et al., 2008; Sharp et al., 2009). As a consequence, we will use used the concept of open-source software as a social good (Huang et al., 2021) as an extreme example (i.e., the developers help solve a social problem, but do not receive a payment) for designing a fourth one payoff function in our experiment.
@@ 3 Study Protocol
As explained previously, our study involves involved two data-collection processes, a survey and a laboratory experiment. In Table 1, we provide an overview of our intended study goals based on the Peer Community In Registered Reports (PCI RR) [^3] https://rr.peercommunityin.org/ study design template, which we will explain in more detail in this section. Our whole study design has received approval from the local Ethics Review Board of the Department for Mathematics and Computer Science at Eindhoven University of Technology, The Netherlands, on October 24, 2022 (reference ERB2022MCS21).
Table 1: PCI RR study design template for our initial study design. In the last column, we explain whether and why we deviated from this design (all changes were approved by the recommender).
question hypothesis sampling plan analysis plan sensitivity rationale interpretation disproved theory deviations Which payoff functions are applied/preferred in SE practice? (survey) N/A At least 30 participants (personal contacts and social media). We will analyze analyzed the absolute frequency of the combinations of payment components. We computed the mean values of the weights for the MA and MP combinations. N/A If MAIT and MPIT should be were identical, we will would have reduced the number of treatments from four to three. N/A We conducted an additional iteration of the (translated) survey with eight participants from a German company to achieve our anticipated sample size. How do different payoff functions impact the performance of participants in SE experiments? (experiment) H1: Participants without performance-based incentivization (NPIT) have on average a worse performance than those with performance-based incentivization (e.g., OSIT, MAIT, MPIT).
@@
We aimed to recruit at least 80 (20 per treatment) computer-science students of the Otto-von-Guericke University Magdeburg. Furthermore, we conducted an a posteriori power analysis to reason on the power of our tests. If their assumptions are were fulfilled, we used parametric tests to compare between the treatments. Otherwise, we will employed non-parametric tests. For H1, we will used pairwise compare comparisons of the performance-independent treatment to the other treatments: NPIT vs. MPIT NPIT vs. MAIT NPIT vs. OSIT For H2, we will used pairwise compare comparisons of the performance-dependent treatments: MPIT vs. MAIT MAIT vs. OSIT OSIT vs. MPIT In total, we will would compute up to six pairwise tests to compare the (at most) four treatments with one another and will corrected for multiple hypotheses testing (Holm-Bonferroni method). We will also conducted regression analyses using the treatments as categorical variables (NPIT as base) and age, gender, experience, as well as arousal as exogenous variables Due to our experimental design, we faced the issue of multiple hypotheses testing. We addressed this issue by applying the Holm-Bonferroni correction. We find support for H1, if our participants’ performance in NPIT is significantly lower than in any other of our experimental treatments at $p<0.05$—after correcting with the Holm-Bonferroni method: (NPIT $<$ MPIT) OR (NPIT $<$ MAIT) OR (NPIT $<$ OSIT). Confirming H1 means that the performance is better in the specific treatment with performance-based incentives compared to NPIT. This implies that if performance plays a role in a software-engineering experiment, performance-based incentivization should be considered.
@@
There is no theory focusing on the role of incentives in software engineering. Incentivization in software-engineering experiments is scarcely applied. Our results could can improve experimental designs in software engineering by guiding researchers when and how to use incentives in their experiments. While we anticipated the possibility that MAIT and MPIT would be identical and should be merged, this did not happen. However, we found that MAIT and NPIT were essentially identical, which is why we merged these two. The changes were made prior to the commencement of the experiment and were approved by PCI RR on 06 Dec 2022.
@@ 3.1 Survey Design
Goal. With our survey, we aimed to explore {enumerate*}[label=)]
which payment components (e.g., wages only, bug bounties) are most applied (MA) in practice and
@@
which payment components are most preferred (MP) by practitioners. We display an overview of these payment components with concrete examples in Table 2. Our intention is was to understand what is actually employed compared to what would be preferred as a payment schema to guide the design of our experiment.
@@
Structure. To achieve our goal, we created an online questionnaire with the following structure (cf. Table 3). At first, we will welcomed our participants, informing them about the survey’s topic, duration, and their right to withdraw from our experiment at any point in time without any disadvantages. Furthermore, we will asked for consent to collect, process, and publish the data in anonymized from form. To allow for questions, we will provided the contact data of at least one author on this the first page. Then, we will asked about each participant’s background to collect control variables, for instance, regarding their demographics, role in their organization, the domain they work in, and experience with code reviews. These background questions allow us to monitor whether we have acquired a broad sample of responses from different organizations, and thus on varying practices. Our goal is was to mitigate any bias caused by external variables, such as the organizations’ culture. Also, we will discarded the answers of participants one participant who had no experience with code reviews. Based on the participants’ roles, the online survey will showed the questions on the payment structures in an adaptive manner. We designed these questions as well as their answering options based on established guidelines and our experiences with empirical studies in software engineering (Siegmund et al., 2014; Nielebock et al., 2019; Krüger et al., 2019).
@@
To explore the payment components (target variables), we will displayed the ones we summarize in Table 2. We will used a checklist in which a participant can could select all components that are applied in their organization. Each selected component will have had a field in which the participant can could enter a percentage to indicate to what extent that component impacts impacted their payment (e.g., 80 % wage and 20 % bug bounty). Then, we presented the same checklist and fields again. This time, the participant shall should define which subset of the components they would prefer to contribute with what share to the payment. While we presented this second list as is to any management role (e.g., project manager, CEO), we asked software engineers (e.g., developer, tester) to decide upon those components from the perspective of being the team or organization lead. To prevent sequence effects, we will randomized the order in which the two treatment questions occured (applied and preferred). Finally, we asked each participant to indicate how many hours per week they worked unpaid overtime—which represents a type of performance penalty for our payoff functions—and allowed them to enter any additional comments on the survey.
@@
Table 2: List of components of payment we will asked about in our survey to design payoff functions for the experiment. Note that the term check refers to participants selecting or deselecting a line of code during our experiment (i.e., marking them as buggy or correct as can be seen in Figure 1).
@@
Sampling Participants. We will invited personal contacts and collaborators from different organizations, involving software developers, project managers, and company managers. Note that we excluded self-employed or freelancer developers who typically ask for a fixed payment for a specific task or project. In addition, we will distributed a second version (to distinguish both populations) of our survey through our social media networks. In consultation with the PCI Recommender (December 6, 2022), we surveyed an additional sample of eight employees from a company to obtain a larger sample size. For this additional sample, we translated the questionnaire into German. We will tested whether there are differences between the samples regarding our variables of interest. If the MA and MP incentives are were identical in both all samples, we will collapse would have collapsed the data. Otherwise, we will build would have built on the sample of our personal contacts only. This allows allowed us to have a higher level of control over the participants’ software-engineering background, and their experience with code reviews.
@@
Our goal is was to acquire at least 30 responses to obtain a reasonable understanding of applied and preferred payments. Since we do did not evaluate the survey data using inferential statistics, we based our sample-size planning on the limited access to a small, specialized number of potential participants. Note that we will did not pay incentives for participating in the survey. We expected that the survey will would take 10 minutes at most, but will and did verify the required time and understandability of the survey through test runs with three PhD students from our work groups.
@@
Analysis Plan. To specify the payoff functions for our experiment, we will considered the absolute frequency of combinations of different payment components. Precisely, to identify the MA and MP combinations, we will choose chose the respective combination that was selected by the largest number of respondents (i.e., modal value). For these two combinations, we will computeed the mean values for their weights. We will performed a graphical-outlier analysis using boxplots Tukey (1977), excluding participants with extreme values (i.e., three inter quartile ranges above the third quartile or below the first quartile). As an example, assume that most of our participants would state to prefer the combination of fixed wages (with a weight of 75 % on average) and bug bounties (25 % on average). Then, we would define a cost function as $0.75\cdot payment_{fix}+0.25\cdot(bugs_{correct}\cdot reward_{quality})$.
@@
Threats to Validity. Our survey reliesd mostly on our personal contacts, which may bias have biased its outcomes. We can mitigated this threat, since we have a broad set of collaborators in different countries and organizations. Moreover, defining the “ideal” payoff function for practitioners may pressure the participants, is hard to define (e.g., considering different countries, organizational cultures, open-source communities, or expectations), and challenging to measure (e.g., what is preferred or efficient). However, this is due to the nature of our experiment and the property we study: financial incentives. Consequently, these threats remain and we have to discuss their potential impact, which can only be mitigated with an appropriately large sample population.
Table 3: List of variables we will check checked in our survey.

### v7

@@ Abstract
In software-engineering research, many empirical studies are conducted with open-source or industry developers. However, in contrast to other research communities like economics or psychology, only few experiments use financial incentives (i.e., paying money) as a strategy to motivate participants’ behavior and reward their performance. The most recent version of the SIGSOFT Empirical Standards mentions payouts only for increasing participation in surveys, but not for mimicking real-world motivations and behavior in experiments. Within this article, we report a controlled experiment in which we tackled this gap by studying how different financial incentivization schemes impact developers. For this purpose, we first conducted a survey on financial incentives used in the real-world, based on which we designed three incentivization schemes: (1) a performance-dependent scheme that employees prefer, (2) a scheme that is performance-independent, and (3) a scheme that mimics open-source development. Then, using a between-subject experimental design, we explored how these three schemes impact participants’ performance. Our findings indicate that the different schemes can impact participants’ performance in software-engineering experiments. Due to the small sample sizes, our results are not statistically significant, but we can still observe clear tendencies. Our results are not statistically significant, possibly due to small sample sizes and the consequent lack of statistical power, but with some notable trends that may inspire future hypothesis generation. Our contributions help understand the impact of financial incentives on participants in experiments as well as real-world scenarios, guiding researchers in designing experiments and organizations in compensating developers.
@@ 1 Motivation
Interestingly, there are no guidelines or recommendations on using financial incentives in software-engineering experimentation. For instance, the current SIGSOFT Empirical Standards [^1] https://github.com/acmsigsoft/EmpiricalStandards (Ralph, 2021), as of January 22, 2024 (commit 9374ea5), mention incentives solely in the context of longitudinal studies and rewarding participation in surveys to increase participation. Also, to the best of our knowledge and based on a literature review, financial incentives that reward participants’ performance during an experiment are not used systematically in empirical software engineering. Although some studies broadly incentivize performance (e.g., Sayagh et al. (2020) or Shargabi et al. (2020)), these do not aim to improve the validity of the experiment, only participation. Furthermore, we know from experimental economics (Charness and Kuhn, 2011; Carpenter and Huet-Vaughn, 2019) that finding a realistic (and thus externally valid) way to reward performance is challenging and no simple one-fits-all solution exists. For instance, the performance of open-source developers depends less on financial rewards than those of industrial developers (Baddoo et al., 2006; Ye and Kishida, 2003; Huang et al., 2021; Beecham et al., 2008).
@@
As a step towards understanding and systematizing the potential of using financial incentives in software engineering experimentation, we have conducted a two-part study comprising a survey and a controlled experiment in the context of bug detection through code reviews (Krüger et al., 2022). First, we used a survey with practitioners to elicit real-world incentivisation schemes on bug finding. In the survey, we distinguished between the schemes most participants prefer and those actually employed. Building on the results, we defined one payoff function for our experiment. Please note that we originally planned to have two functions from the survey, one for the most applied (MA) and one for the most preferred (MP) incentives (Krüger et al., 2022). However, the survey responses for the MA incentives were identical to no performance-based incentives, which we added as a control treatment anyway. To extend our experiment, we added two more payoff functions: one that is performance-independent and one that resembles the motives of open-source developers. We derived the latter function using the induced-value method established in experimental economics (Smith, 1976; Weimann and Brosig-Koch, 2019), which induces a controlled willingness of participants to achieve a desired goal (i.e., identify a bug) or obtain a certain good during an experiment by mimicking its monetary value (e.g., a reward). Second, we employed our actual between-subject experiment to explore to what extent each of the three payoff functions impacts the participants’ behavior. Overall, we primarily contribute to improving researchers’ understanding of whether and how financial incentives can help software engineering experimentation. However, our experiment can also help reveal whether different incentivisation schemes could improve practitioners’ motivation. Our survey and experimental design artifacts are available for peer-reviewing.
@@
    We share our artifacts, including the design and results of our survey as well as experiment in anonymous form within a persistent open-access repository. [^2] https://osf.io/mcxed/?view_only=602088776ce5498597c473e74bbe0394
@@ 2 Related Work
Experiments in software engineering are comparable to “real-effort experiments” in experimental economics, which involve participants who solve certain tasks to increase their payoffs. Consequently, we built on experiences from the field of experimental economics, which involves a large amount of literature on how and when to use financial incentives in real-effort experiments (van Dijk et al., 2001; Greiner et al., 2011; Gill and Prowse, 2012; Erkal et al., 2018). For instance, some findings indicate gender differences regarding the impact of incentivization schemes, which we have to consider during our experiment. In detail, research has shown that men choose more competitive schemes (e.g., tournaments, performance payments). Similarly, participants with higher social preferences select such competitive schemes more rarely (Niederle and Vesterlund, 2007; Dohmen and Falk, 2011). We considered such factors when analyzing the results of our experiment (e.g., comparing gender differences if the number of participants allows).
@@
Software-engineering researchers have investigated the motivations of open-source developers to a much greater extent (Gerosa et al., 2021; Hertel et al., 2003; Hars and Ou, 2002; Ye and Kishida, 2003; Huang et al., 2021). From the economics perspective, open-source systems represent a public good (Bitzer et al., 2007; Lerner and Tirole, 2003): they are available to everyone and their consumption do not yield disadvantages to anyone else. A typical problem of public goods is that individual and group incentives collide, which usually leads to an insufficient provision of the good. While typical explanations for open-source development focus on high intrinsic motivation to contribute or learn, this is not always the case. For instance, Roberts et al. (2006) show that financial incentives can actually improve open-source developers’ motivation (in terms of contributions). Still, financial incentives are at least not always the predominant motivators for software developers (Beecham et al., 2008; Sharp et al., 2009). As a consequence, we used the concept of open-source software as a social good (Huang et al., 2021) as an extreme example (i.e., the developers help solve a social problem, but do not receive a payment) for designing one payoff function in our experiment.
@@ 3 Study Protocol
As explained previously, our study involved two data-collection processes, a survey and a laboratory experiment. In Table 1, we provide an overview of our intended study goals based on the Peer Community In Registered Reports (PCI RR) [^3] https://rr.peercommunityin.org/ study design template, which we explain in more detail in this section. Our study design was based on guidelines for using financial incentives in software-engineering experimentation (Krüger et al., 2024) and has received approval from the local Ethics Review Board of the Department for Mathematics and Computer Science at Eindhoven University of Technology, The Netherlands, on October 24, 2022 (reference ERB2022MCS21).
Table 1: PCI RR study design template for our initial study design. In the column deviations, we explain whether and why we deviated from this design (all changes were approved by the recommender).
question hypothesis sampling plan analysis plan sensitivity rationale interpretation disproved theory deviations observed outcome Which payoff functions are applied/preferred in SE practice? (survey) N/A At least 30 participants (personal contacts and social media). We analyzed the absolute frequency of the combinations of payment components. We computed the mean values of the weights for the MA and MP combinations. N/A If MAIT and MPIT were identical, we would have reduced the number of treatments from four to three. N/A We conducted an additional iteration of the (translated) survey with eight participants from a German company to achieve our anticipated sample size. The most commonly applied payments are fixed. The most commonly preferred one is a combination of fixed payment and company-performance-dependent bonus. How do different payoff functions impact the performance of participants in SE experiments? (experiment) H1: Participants without performance-based incentivization (NPIT) have on average a worse performance than those with performance-based incentivization (e.g., OSIT, MAIT, MPIT).
@@
We aimed to recruit at least 80 (20 per treatment) computer-science students of the Otto-von-Guericke University Magdeburg. Furthermore, we conducted an a posteriori power analysis to reason on the power of our tests. If their assumptions were fulfilled, we used parametric tests to compare between the treatments. Otherwise, we employed non-parametric tests. For H1, we used pairwise comparisons of the performance-independent treatment to the other treatments: NPIT vs. MPIT NPIT vs. MAIT NPIT vs. OSIT For H2, we used pairwise comparisons of the performance-dependent treatments: MPIT vs. MAIT MAIT vs. OSIT OSIT vs. MPIT In total, we would compute up to six pairwise tests to compare the (at most) four treatments with one another and corrected for multiple hypotheses testing (Holm-Bonferroni method). We also conducted regression analyses using the treatments as categorical variables (NPIT as base) and age, gender, experience, as well as arousal as exogenous variables Due to our experimental design, we faced the issue of multiple hypotheses testing. We addressed this issue by applying the Holm-Bonferroni correction. We find support for H1, if our participants’ performance in NPIT is significantly lower than in any other of our experimental treatments at $p<0.05$—after correcting with the Holm-Bonferroni method: (NPIT $<$ MPIT) OR (NPIT $<$ MAIT) OR (NPIT $<$ OSIT). Confirming H1 means that the performance is better in the specific treatment with performance-based incentives compared to NPIT. This implies that if performance plays a role in a software-engineering experiment, performance-based incentivization should be considered.
@@
There is no theory focusing on the role of incentives in software engineering. Incentivization in software-engineering experiments is scarcely applied. Our results can improve experimental designs in software engineering by guiding researchers when and how to use incentives in their experiments. While we anticipated the possibility that MAIT and MPIT would be identical and should be merged, this did not happen. However, we found that MAIT and NPIT were essentially identical, which is why we merged these two. The changes were made prior to the commencement of the experiment and were approved by PCI RR on 06 Dec 2022. The results of the pre-registered tests were non-significant. Yet, they indicate notable differences that guided our exploratory analysis.
@@ 3.1 Survey Design
Goal. With our survey, we aimed to explore i) which payment components (e.g., wages only, bug bounties) are most applied (MA) in practice and ii) which payment components are most preferred (MP) by practitioners. We display an overview of these payment components with concrete examples in Table 2. Our intention was to understand what is actually employed compared to what would be preferred as a payment schema to guide the design of our experiment.
@@
Structure. To achieve our goal, we created an online questionnaire with the following structure (cf. Table 3). At first, we welcomed our participants, informing them about the survey’s topic, duration, and their right to withdraw from our experiment at any point in time without any disadvantages. Furthermore, we asked for consent to collect, process, and publish the data in anonymized form. To allow for questions, we provided the contact data of one author on the first page. Then, we asked about each participant’s background to collect control variables, for instance, regarding their demographics, role in their organization, the domain they work in, and experience with code reviews. These background questions allow us to monitor whether we have acquired a broad sample of responses from different organizations, and thus on varying practices. Our goal was to mitigate any bias caused by external variables, such as the organizations’ culture. Also, we discarded the answers of one participant who had no experience with code reviews. Based on the participants’ roles, the online survey showed the questions on the payment structures in an adaptive manner. We designed these questions as well as their answering options based on established guidelines and our experiences with empirical studies in software engineering (Siegmund et al., 2014; Nielebock et al., 2019; Krüger et al., 2019).
@@
To explore the payment components (target variables), we displayed the ones we summarize in Table 2. We used a checklist in which a participant could select all components that are applied in their organization. Each selected component had a field in which the participant could enter a percentage to indicate to what extent that component impacted their payment (e.g., 80 % wage and 20 % bug bounty). Then, we presented the same checklist and fields again. This time, the participant should define which subset of the components they would prefer to contribute with what share to the payment. While we presented this second list as is to any management role (e.g., project manager, CEO), we asked software engineers (e.g., developer, tester) to decide upon those components from the perspective of being the team or organization lead. To prevent sequence effects, we randomized the order in which the two treatment questions occured (applied and preferred). Finally, we asked each participant to indicate how many hours per week they worked unpaid overtime—which represents a type of performance penalty for our payoff functions—and allowed them to enter any additional comments on the survey.
@@
Table 2: List of components of payment we asked about in our survey to design payoff functions for the experiment. Note that the term check refers to participants selecting or deselecting a line of code during our experiment (i.e., marking them as buggy or correct as can be seen in Figure 1).
@@
Sampling Participants. We invited personal contacts and collaborators from different organizations, involving software developers, project managers, and company managers. Note that we excluded self-employed or freelancer developers who typically ask for a fixed payment for a specific task or project. In addition, we distributed a second version (to distinguish both populations) of our survey through our social media networks. In consultation with the PCI Recommender (December 6, 2022), we surveyed an additional sample of eight employees from a company to obtain a larger sample size. For this additional sample, we translated the questionnaire into German. We tested whether there are differences between the samples regarding our variables of interest. If the MA and MP incentives were identical in all samples, we would have collapsed the data. Otherwise, we would have built on the sample of our personal contacts only. This allowed us to have a higher level of control over the participants’ software-engineering background, and their experience with code reviews.
@@
Our goal was to acquire at least 30 responses to obtain a reasonable understanding of applied and preferred payments. Since we did not evaluate the survey data using inferential statistics, we based our sample-size planning on the limited access to a small, specialized number of potential participants. Note that we did not pay incentives for participating in the survey. We expected that the survey would take 10 minutes at most, and did verify the required time and understandability of the survey through test runs with three PhD students from our work groups.
@@
Analysis Plan. To specify the payoff functions for our experiment, we considered the absolute frequency of combinations of different payment components. Precisely, to identify the MA and MP combinations, we chose the respective combination that was selected by the largest number of respondents (i.e., modal value). For these two combinations, we computeed the mean values for their weights. We performed a graphical-outlier analysis using boxplots Tukey (1977), excluding participants with extreme values (i.e., three inter quartile ranges above the third quartile or below the first quartile). As an example, assume that most of our participants would state to prefer the combination of fixed wages (with a weight of 75 % on average) and bug bounties (25 % on average). Then, we would define a cost function as $0.75\cdot payment_{fix}+0.25\cdot(bugs_{correct}\cdot reward_{quality})$.
@@
Threats to Validity. Our survey relied mostly on our personal contacts, which may have biased its outcomes. We mitigated this threat, since we have a broad set of collaborators in different countries and organizations. Moreover, defining the “ideal” payoff function for practitioners may pressure the participants, is hard to define (e.g., considering different countries, organizational cultures, open-source communities, or expectations), and challenging to measure (e.g., what is preferred or efficient). However, this is due to the nature of our experiment and the property we study: financial incentives. Consequently, these threats remain and we discuss their potential impact, which can only be mitigated with an appropriately large sample population.
Table 3: List of variables we checked in our survey.